Flippy:

volue of x in 20+5x-10. So, I already know that 20-5x=15 and that 20-(-10)=30 but I don't know what to do after that, do I divide 5 by 30 and get the Value of X?

Flippy:

I spelled it wrong-

lyds:

1). Multiply and Simplify. a) −5(2x-4). = -10x + 20.

candycove:

x = -2. This is assuming you meant = instead of - in the equation. 20 + 5x = 10 -20 -20 5x = -10 /5 /5 x = -2 ✓

candycove:

Or, if you didn't mean = then, x = 6.